UT Mar 13 2022 - Possible DEC time




Night Info

  • Sunset 18:25
  • Nautical Twilight Ends 19:19
  • Astronomical Twilight Ends 19:47
  • Handover from AZ to DEC time 22hr
  • Handover from DEC to AZ time 04hr
  • Moon sets 04:06
  • Astronomical Twilight Starts 05:11
  • Nautical Twilight Starts 05:39
  • Sunrise 06:32

Staff

  • OSA: S. Allanson
  • ISp: J.Schuh
  • Observer: A.Becker
  • AO: J.Power

Night Plan

  • AZ science handover at 22hr, return at 04hr

If conditions are: seeing < 1.2" and no/thin clouds

  • DONE iLocator needs AO quality conditions - 1hr (22:15-23:15)
    • ADC testing
  • Reconfigure for bino-LUCI (23:15-23:30)
  • AO testing 23:30-02:45
    • OT files: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM
      • AO OT Version 78
      • ESM OT Version 36
    • DONE OSCO w/NCPA 23:30-01hr
    • DONE Re-point test 01-01:30hr
      • Confirm patch works
      • Verify it works - only need to acquire target on CCD47 - no need to close the loop
    • Bin 3 IQ test 01:30-02:30
      • Targets selected from OT based on visibility and PA
      • DONE Faint
      • DONE Bright
    • Close Separtion Targets test (0.1"-2") 02:30-03:45
      • in OT: AO STARS - FULL DITHERS *NEW*-->Binaries
        • DONE Target 1 HIP 65069 R=6.9 0.12" separation - need 64x2048 subframe
        • Target 2 BD+00 2888 R=10.7 - separation 0.225" - need 256x2048 subframe
        • Target 5 GD 135 R=14.6 - separation 0.3" - no subframe
        • Target 4 SBSS 1148+564 R=14.9 separation 0.6" - no subframe
        • Target 6 2MASSJ1128+3136 R=13.5 - 1" separation - no subframe
        • Target 7 2MASSJ13061537+2043444 R=8.5 - 1.6" separation -64x2048 subframe
          • Note: given separation and lack of GS only 1 PA script is available - companion should be visible directly below main star
        • Target 3 BD+33 2137 R=10.3 - 2.4" separation no subframe
    • If time permits, and DIMM < 1" - obtain another filter (FeII or Pa-Beta) on Cat's Eye Nebula - visible 01hr-twilight
  • Reconfigure for AZ science 03:45 - if LBC is to be used, reconfigure at 03:30hr

If conditions are: seeing >1.2" and clouds

  • Re-point TCS test (high priority)
    • TCS patch is in place - need to verify its working for SOUL/FLAO on SX
    • verify that re-point works - seeing up to 2"
  • ESM testing in poor conditons - collect FWHMs of stars in ESM on L1 and SL on L2
    • Mixed-mode are TargetName _ESM_SL.xml
    • M67 Twilight-02hr
      • /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/ESM/M67/
      • Bin 1/ R=8.9 (pick any PA)
      • Bin 4/R=16.51 (pick any PA)
    • M92 /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/ESM/M92
      • Bin 1 (R9pt4 scripts) or Bin 4 (R15pt1)
    • LUCI Photometric Standards (any seeing, if so bad can't guide, use track mode) - requires clear skies
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
      • Select stars based on visibility, need to get stars at two different airmasses
      • LUCI is selected as its most likely to be used before handing over from AZ science
    • Poor Weather Science Backup - any non-iLocator good enough conditions qualify
      • Rothberg - Local Mergers
        • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V42.xml
        • NGC 4004 - 21hr-twilight
        • NGC 3597 - 23hr-02hr
        • UGC 9829 - 01:30hr-twilight
        • UGC 9829 - 01:30hr-twilight
        • UGC 10607 - 02hr -twilight
        • UGC 10675 - 02hr-twilight

Night Log

OSCO BD+70 622

  • R=7.48, WFSMag = 8.7
  • Several issues, rolled edges, shell RIPD and ASM went into failure
    • some concern that having LBTI and LUCI WFS's on concurrently caused an issue - remains to be checked
  • After restart - was able to close the loop OK - however a "bite" was seen - J.Power and D.Miller took TN and recorded a movies of it

Repoint Test

  • Test failed - there appears to be a bug where the repoint does not work from the FLAO GUIs, LUCI scripts and IDL scripts don't seem to apply it either via the RunAO.
  • Bayside stages moved rather than the telescope moving
  • Will need to find the bug (SW group action item)

Bin 3 Faint Test KUV 08526+3750

  • UT 07:40 - RA08hr-09hr Star #15 R=14.9, WFSmag=16.1, guider =0.7"
    • IQ rather poor, star is elongated, not many corrections being applied
    • Paused script and changing gains manually to improve IQ
    • N30 Fieldstop is off, will need to fix this later
    • FWHM ~0.09"
  • UT 07:46 - No obvious improvement in corrections via manual gains - even as tip/tilt is ramped up
  • after changes FWHM ~ 0.07" looks elongated - L1 #0126
  • UT 07:56 - optical gain on L1 #0127 - looks PFC jump
  • UT 07:59 - offsetting +1" +1" DETXY to move off bad pixel on LUCI-1
  • UT 08:00 - L1#0128 - NCPA on FWHM~0.078" looks rounder - again looks like a few pixel shift
  • UT 08:06 - resending script with Rmag=16.6 to force to Bin4 (will skip checkflux)
  • UT 08:13 - taking data - L1#0129 - FWHM ~0.1"
  • L1 0121-0130

Bin 3 Bright Target G58-40

  • RA10hr-11hr folder Star #10 PA=180
  • Rmag=13.92, WFSMag=13.5 (that's surprising!)
  • Ended brighter than expected so moving to another target

Bin 3 Bright Target PG 1325+102

  • PA=180, 12hr-13hr folder Star #7 Rmag=14.2, WFSMag=15.9
  • unexpectedly too faint, despite being 80deg from Moon

Bin 3 Bright Target PM J15220+5901

  • PA=180,RA14-15hr folder Star#6 Rmag=13.5, WFSMag=13.9
  • 80 deg from Moon
  • noise on OCAM again
  • damnit....undershot

Bin 3 Bright Target EGGR 440

  • PA=120, RA14hr-15hr folder Star#7 Rmag=14.2, WFSMag=15.5, instrument rotator=167
  • Spectroscopic Binary
  • Freq=750Hz - so ACE table thinks this is at the bright end
  • Frames in cube show varying IQ L1 #0132
  • Also looks unstable in L1 #0134 - is seeing that variable every 5sec?
  • NCPA does improve the IQ quite a bit, can even see airy rings, within cube there's still instability in the IQ from frame to frame
  • L1 0132-0142 - NCPA off
  • L1 0143- NCPA On

Binary Separation Test

Test#1 HIP 65069 R=9 0.12" separation
  • 64x2048 subframe Rmag=6.8, WFSMag=8.8
  • NCPA was on
  • Had to fix scripts (B.Rothberg error) - so a few minutes downtime
  • UT 09:57 - Safeskipping in an offset - should be paused
  • L1 0164 - loop paused - image is bad - J.Power checking on what happened
  • UT 09:59 - safeskipping, bite in mirror GUI
  • UT 10:00 - optical loop data being taken - no LUCI images being taken at the moment
  • UT 10:01 - 22%, now 40% safeskipping - and crash......loop opened
  • L1 0149-0164 NCPA ON
  • UT 10:05 - running again (changed script to have smaller offsets) - without NCPA to see if get same instability
  • UT 10:09 - subwindow needs to be fixed
  • UT 10:17 - so far no safeskippiny
  • - there was confusion about the target returning from an offset 0.5" from the cetner of the frame - it was centered up on an offset not 0,0
  • UT 10:29 - back to 0,0 - just sitting here with NCPA on loop closed, looking for loop instabilities
  • CONCLUSION SO FAR: NCPA may build up instability over time with close seprations - next test - NCPA on and NCPA off for close separation tests.
  • L1 0165-0186

UT 10:31 - handing over to AZ to reconfigure for LBCs

-- BarryRothberg - 12 Mar 2022
Topic revision: r3 - 13 Mar 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback